OpenStudy (purexyz):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (anonymous):

Since, parallel lines have same slope therefore, Slope of required line = slope of 2x + 3y = 5 i.e. 3y= -2x+5 i.e. \[\huge y= \frac{-2}{3} x + 5 \] Comparing the above equation with \[\huge y= m x +c\] We find slope m= \[\huge m=\frac{-2}{3} \] Silnce line passes through the point (9, -3) Hence, eq of line in slope- point form is given as; \[\huge y-y_1= m(x-x_1)\] Where x1 = 9 & y1= -3 substituting the values of m, x1 & y1 in the above equation we find: \[\huge y-(-3)= \frac{-2}{3}(x-9)\] \[\huge y+3= \frac{-2}{3}(x-9) \rightarrow \huge 3( y+3)= -2(x-9)\] \[\rightarrow \huge 3y+9= -2x+18\] \[\rightarrow \huge 2x+3y=18+9\] \[ \huge \rightarrow 2x+3y=27\] hence \[ \huge 2x+3y=27 \] is the required equation of the line. @purexyz

OpenStudy (anonymous):

@purexyz

OpenStudy (purexyz):

thanks